OpenStudy (superhelp101):

Select the best possible first step to solving the system by first eliminating the x variable. x - 3y = 12 2x + 5y = 2 Multiply the first equation by 2. Multiply the second equation by 2. Multiply the first equation by -2. Multiply the second equation by -2. Doesn't a and c work?

OpenStudy (superhelp101):

@amistre64

OpenStudy (superhelp101):

can you help me?

OpenStudy (anonymous):

multiply the 1st eq by 2

OpenStudy (superhelp101):

2x-6y=24 2x+5y=2 then you add them

OpenStudy (muzzack):

i would say A

OpenStudy (anonymous):

if you multiply it by 2 then subtract them if you multiply it by -2 then add them up.

OpenStudy (superhelp101):

yes but that's were I was confused.

OpenStudy (superhelp101):

both eliminate x in a different way

OpenStudy (superhelp101):

right :/

OpenStudy (anonymous):

yes...

OpenStudy (mosaic):

True. But multiplying by -2 and adding is a bit simpler than subtracting.

OpenStudy (superhelp101):

but i can't do 2 answers

OpenStudy (anonymous):

yuuuppp...but use which you think is easeir...

OpenStudy (superhelp101):

ok thx

OpenStudy (superhelp101):

@yungrichneegu @radar what do you think?

OpenStudy (superhelp101):

@mathmate

OpenStudy (texaschic101):

I would multiply the first equation by -2 and then add them....that will eliminate the x's but your right...you could multiply the first equation by 2, and then subtract them. Either way will work.

OpenStudy (superhelp101):

Alright, thanks! so if I pick either one its right?

OpenStudy (texaschic101):

I, myself, think the adding is simplier...but you can choose either way. x - 3y = 12 --->(-2) 2x + 5y = 2 --------------- -2x + 6y = -24 (result of multiplying by -2) 2x + 5y = 2 ---------------add 11y = - 22 y = -2 or x - 3y = 12 --->(2) 2x + 5y = 2 ------------- 2x - 6y = 24 -(2x + 5y = 2) ------------- now just distribute the negative through the parenthesis so now we have 2x - 6y = 24 -2x - 5y = -2 --------------- -11y = 22 y = -2

OpenStudy (superhelp101):

okay!

OpenStudy (mathmate):

Personally, I suggest multiplying by a positive number and decide later to add or subtract. This will avoid multiplying by a negative number and get lost with the sign. But this is a subjective opinion. Everyone has a different favourite procedure.